This problem is about a fast-food business. As it turns out, 60% of their business is via the drive-through window, and 40% is counter-service. Data from the registers reveals that 70% of the transactions at the drive-through are via credit/debit cards, while 80% of the transactions at the counter are via credit/debit cards. (The other transactions are via cash.) What is the probability....A.that a random transaction uses cash? B.that a random transaction uses credit cards? C.that a transaction was a credit-card transaction and at the drive-through window? D.that a transaction was a cash transaction and was at the drive-through window? E.that a random transaction was at the counter? F.that a random transaction was at the drive-through window?
1
Expert's answer
2020-12-02T19:22:17-0500

What is the probability....

A.that a random transaction uses cash?

P = 0.6*(1 - 0.7) + 0.4*(1 - 0.8) = 0.18 + 0.08 = 0.26

B.that a random transaction uses credit cards?

P = 1 - 0.26 = 0.74

C.that a transaction was a credit-card transaction and at the drive-through window?

P = 0.7*0.6 = 0.42

D.that a transaction was a cash transaction and was at the drive-through window?

P = (1 - 0.7) * 0.6 = 0.18

E.that a random transaction was at the counter?

P = 0.4

F.that a random transaction was at the drive-through window?

P = 0.6